OpenStudy (thickgirl_angel):

An anthropologist finds bone that her instruments measure it as 0.012% of the amount of Carbon-14 the bones would have contained when the person was alive. How long ago did the person die? The half life of carbon 14 is 5,730 years. Round your answer to the nearest thousand.

OpenStudy (kgrendel0324):

\[(1/2)^{\frac{ t }{ 5730 }}=0.005\]

OpenStudy (kgrendel0324):

solve for t

OpenStudy (kgrendel0324):

\[\frac{ t }{5730}=\frac{ \ln(.005) }{\ln(.5)}\] \[t=\frac{ 5730\ln(.005) }{\ln(.5) }\]

OpenStudy (kgrendel0324):

\[5730*\frac{ \log(0.005) }{ \log(0.5) }\]

OpenStudy (kgrendel0324):

43,799 years Rounded to the nearest thousand: 44,000